[Hallc_running] [New Logentry] RC Report - 8/29/22

arunts at jlab.org arunts at jlab.org
Mon Aug 29 09:55:02 EDT 2022

Logentry Text:
We moved to the SHMS 10 deg configuration (page 42 of the run plan) during beam recovery yesterday and took data at 38 uA on LD2 and 40 uA on dummy target. If we have the same rate of beam trips as today, and no major outages, we will reach this goal around 20:00 Monday swing (Aug 29). We plan to move to either the 7.5 deg or the 6 deg configuration later today which can be discussed at the RC meeting. Speaking of which: 

Let's have a brief RC meeting today at 11:00 AM EST.

Meeting ID: 160 710 0885 - Passcode: 595071


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/4035316
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20220829/13ad7b5e/attachment.html>

More information about the Hallc_running mailing list